OpenStudy (anonymous):

log2 1/64 (the 2 is slightly below the log)

OpenStudy (turingtest):

instead of writing "the 2 is slightly below the log" you can write "log base 2" which is the proper terminology

OpenStudy (anonymous):

thankx

OpenStudy (anonymous):

2^x = 1/64 x= -6